# Universal Jailbreak Suffixes Are Strong Attention Hijackers

### v1

@@ 5.1 Formalizing Hijacking
Following Elhage et al. (2022), for transformer-based LMs Vaswani et al. (2017), the representation of token in layer is given by:
where denotes the MLP sub-layer, and is the attention sub-layer. Crucially, only the latter incorporates information from previous tokens. Following Kobayashi et al. (2020, 2021), this attention term decomposes as a sum of the transformed vectors:
@@
Next, we build on previous approaches to quantify the contribution of model components Kobayashi et al. (2021); Mickus et al. (2022), and define a dot-product-based dominance metric to assess the contribution of a token subsequence , in a given direction , for a specific layer :
Dominance score. To quantify the contributors to chat’s contextualization, we evaluate the attention sub-layer output in layer , by setting , . in Eq. 2. can be any token subsequence preceding chat[-1] (e.g., adv).222 Note that summing over the contributions of all tokens yields (i.e., ). To simplify the analysis to follow, we select the last token position before generation (chat[-1]) as a single representative token from chat. Formally:
Intuitively, this metric measures how much influences a target subsequence (chat[-1]), thus we refer to it as ’s dominance score. Mathematically, it captures the magnitude of ’s contribution in the direction of the total attention output at layer , which is in turn added to the residual stream (). When has a markedly higher dominance score than other token subsequences’, we say it hijacks the context.
@@ 5.2 Hijacks
We use the dominance score (Eq. 5.1) to analyze how different subsequences contribute to the jailbreak-critical chat tokens across various prompt distributions. We find that GCG suffixes aggressively hijack chat and strongly suppress instr’s dominance, far more than other prompts.
We first compare two prompts with the same instruction—one using a random suffix (“!!..!”, failed sample) and the other a GCG suffix (successful)—and visualize their dominance scores across layers in Fig. 5. Unlike the random suffix, GCG dominates the chat representation from later layers (), practically hijacking the contextualization process, while the instr’s influence gradually vanishes already from early layers (). Observing that this gap concentrates at the upper layers, we next analyze these.
Experimental setting. We extend our analysis by averaging the dominance score across the upper half of the layers (which can be seen as accumulating the respective area in Fig. 5) and evaluating it on 30 harmful instructions, each paired with suffixes from six distributions (examples in Tab. 3):
@@
Results. Analyzing the suffixes’ dominance scores (Fig. 6), we find that GCG suffixes highly suppress the instruction, while exhibiting irregular dominance in contextualization, contributing a magnitude of over compared to other suffix distributions, including handcrafted jailbreaks. We observe similar trends aggregating on all the layers (Fig. 11, App. B.3). Furthermore, GCG’s hijacking remains unusual even compared to dominant suffixes in benign instructions (from AlpacaEval; Fig. 12, App. B.3), underscoring the distinctiveness of this mechanism.
In general, successful GCG jailbreaks require heavily suppressing instr’s contribution, with adv strongly hijacking chat by contributing a large magnitude to the former’s representation (Figs. 6(a)–6(b)); as seen in §4, this predominant and unique mechanism is also critical for jailbreaks. GCG’s dominance extends to the entire prompt (Fig. 6(c)), suggesting it additionally suppresses the influence of template tokens in pre-chat (e.g., <bos>) and chat itself—a pattern also visible in Fig. 5. Moreover, handcrafted jailbreaks’ relatively strong hijacking may explain their effective use as suffix initializers in jailbreak optimizers Liu et al. (2024).
@@ 6 Hijacking is Key for GCG Universality
Experimental setting. We next evaluate Gemma2, and in App. B.4 show similar results for Qwen2.5. We sample 350 GCG suffixes of diverse universality, along with 30 harmful instructions. For each suffix, we average its adv dominance score in layer across the different instructions, referring to this measure as the suffix’s hijacking strength, and comparing it to the suffix’s universality score (calculated on a larger set of instructions; §3.1). Notably, this calculation involves a single forward pass.
Results. Fig. 7 shows the Spearman correlation between universality and layer-wise hijacking strengths (including a summation over all the layers), using an instruction set (of size 10) disjoint from the evaluations to follow. We find that dominance in the initial and final layers does not correlate with universality, whereas hijacking strength in later-mid layers does. Notably, layers 18–21 yield the highest Spearman correlations; specifically, layer 20 achieves a moderate correlation Schober et al. (2018) of , -value , and confidence interval of Fieller et al. (1957), with similar values for the other 20 instructions used in further evaluation.
@@
Notably, similar trends hold when measuring hijacking strength using a much smaller instruction set, even a single random harmful instruction (i.e., simply comparing suffixes’ dominance scores under an instruction; Fig. 8(b)). Additionally, to control for the possible effect of jailbreak success, we repeat this analysis, using only failed attack samples (in particular, instruction-suffix pairs that elicit model refusal). As shown in Fig. 8(c), the trend persists, indicating the internal hijacking mechanism varies across suffixes, even when all produce the same refusal outcome.
Lastly, we demonstrate that results and correlations repeat with other variants of hijacking strength (Fig. 13, App. B.4), underscoring the role of the attention scores, and that the hijacking can be inspected along a few directions. Specifically, hijacking strength (in layer 20) can also be computed as: (i) a statistic aggregating top advchat attention scores (which directly scale the transformed vectors, and could control the hijacking strength; Eq. 1); (ii) dominance score w.r.t. a principal direction in GCG jailbreaks (replacing the sample-specific attention activations in Eq. 5.1), derived with difference-in-means on failed and successful GCG samples Arditi et al. (2024).
@@ 7 Practical Implications
Here we translate our insights on GCG jailbreak into two practical strategies. First, we boost GCG’s universality via a hijacking-enhanced objective (§7.1). Then, we mitigate GCG jailbreaks by suppressing hijacking during inference (§7.2).
@@ 7.1 Boosting GCG Universality With Hijacking Enhancement
We leverage our insights on the relationship between hijacking and universality (§6), to encourage hijacking during GCG’s optimization. We show it is possible to use this method to obtain universal suffixes with a reduced computational cost.
@@
Existing approaches. Universal suffix-based jailbreaks are typically crafted by running the original affirmation objective on multiple harmful instructions (GCG-Mult; Zou et al. (2023b)). However, optimizing GCG across instructions significantly increases computational cost—matching that of separate single-instruction runs.
@@
Hijacking-based approach (GCG-Hij). We propose a modified objective that is optimized against a single instruction, thus preserving the computational efficiency of single-instruction GCG. Specifically, motivated by the fact that attention scores scale transformed vectors magnitude (thus enhance hijacking; Eq. 1), and increase with universality in middle layers (§6), we define an attention-score-based proxy objective (), which is then added to GCG’s affirmation loss to form GCG-Hij’s loss ():
@@
Experimental setting. We select 10 random instructions for attacks (disjoint from evaluation), run GCG and GCG-Hij on each instruction separately, and optimize GCG-Mult on all 10 instructions simultaneously.333 We select values by line search on a dev set (disjoint from evaluation): , , and for Gemma2, Qwen2.5, and Llama3.1, respectively, generally finding effective. For layers, we set and . We repeat this for 3 random seeds. To test whether GCG-Hij is more likely to yield universal suffixes from single-instruction optimization, we compute the average universality (§3.1) on the 10 suffixes crafted with GCG and GCG-Hij (Avg. Univ.). Then, under a unified budget of 10 optimized instructions, we compare GCG and GCG-Hij to GCG-Mult, reporting whether either surpasses GCG-Mult (Win Over GCG-Mult) and the fraction of such wins (% of suffixes won).
GCG-Hij boosts universality. Results are shown in Tab. 1. First, GCG-Hij achieves superior average universality, compared to GCG (1.1–5); that is, GCG-Hij is more likely to create a more universal suffix, at a cost as low as the original single-instruction GCG. Second, for the same computational budget as GCG-Mult’s, GCG-Hij consistently produces (1) universal suffixes that surpass the single suffix crafted by the former. App. B.5 includes the testing of another GCG variant that replaces the default initialization with handcrafted jailbreaks (which exhibit strong hijacking; §5), without modifying the objective. This variant mostly yields universality boosts similar to GCG-Hij’s (Fig. 16), while giving the attack a strong head start in hijacking (Fig. 17).
7.2 Mitigating GCG Jailbreak with Hijacking Suppression
@@
Through our analysis, we found that GCG adversarial suffixes (adv) hijack chat’s representation in an irregular and often extreme manner, particularly for universal suffixes (§5–6), and that this hijacking underlies attack effectiveness (§4). We therefore hypothesize that surgically suppressing this hijacking could defend against GCG jailbreaks with minimal effect on benign prompts. To test this, we introduce and evaluate a training-free framework for Hijacking Suppression.
@@
Starting with (a), we consider as candidates all transformed vectors departing from the user input tokens (excluding special tokens) to chat tokens, denoted inputchat. Notably, for GCG prompts, suppressing a large portion of these vectors (i.e., advchat, see §4.1), eliminates the attack. We use this general superset (inputchat) so the framework remains applicable to any prompt, including benign ones, without prior prompt knowledge.
@@
Next, for (b), we assign each vector in the superset a score, and select the top-1%. Specifically, we use the attention score () for each transformed vector (), as it mathematically scales the vector (Eq. 1) potentially amplifying hijacking strength, and empirically, higher top-1% scores correlate with GCG suffix universality (§6). While we prioritize simplicity, other scoring methods may better disentangle jailbreak from model utility, which we defer to future work to explore.
@@
Finally, for (c), we suppress the top 1% of transformed vectors by scaling their magnitude by : 444The transformed vector update is applied before layer normalization and is equivalent to reducing the corresponding post-softmax attention score.
@@
Experimental setting. We apply Hij. Suppr. with on different models,555We found balances robustness and utility; further tuning may improve results. and evaluate the effect on: (i) model robustness, by measuring the attack success rate on a challenging custom dataset of 1.5K GCG jailbreak prompts, using harmful instructions from AdvBench Zou et al. (2023b) and StrongReject Souly et al. (2024), each appended to various GCG suffixes, that originally led to diverse attack success; and on (ii) model utility, using AlpacaEval Li et al. (2023) and MMLU Hendrycks et al. (2021). See App. A.3 for more technical details.
| Attack Success () | Utility () | |||
| GCG | AlpacaEval | MMLU | ||
| Gemma2 | Initial | |||
| +Hij. Suppr. | ||||
| Qwen2.5 | Initial | |||
| +Hij. Suppr. | ||||
| Llama3.1 | Initial | |||
| +Hij. Suppr. | ||||
@@
Hij. Suppr. improves robustness vs. GCG. Tab. 2 presents Hij. Suppr.’s effect on attack success. GCG’s attack success is largely eliminated, being reduced by a factor of –. On the other hand, we spot a consistent yet slight drop in model utility of 2% decrease on MMLU and AlpacaEval. AlpacaEval responses remain highly similar after applying Hij. Suppr., with average RougeL scores of –, indicating minimal change Lin (2004). Still, we expect further refinement of the framework (e.g., the scoring step, (b)) to improve robustness-utility tradeoff.

### v2

@@ 5.1 Formalizing Hijacking
Following elhage2022solu, for transformer-based LMs vaswani2017attention, the representation of token in layer is given by:
where denotes the MLP sub-layer, and is the attention sub-layer. Crucially, only the latter incorporates information from previous tokens. Following kobayashiAttentionNotOnly2020a; kobayashiIncorporatingResidualNormalization2021, this attention term decomposes as a sum of the transformed vectors:
@@
Next, we build on previous approaches to quantify the contribution of model components kobayashiIncorporatingResidualNormalization2021; mickusHowDissectMuppet2022, and define a dot-product-based dominance metric to assess the contribution of a token subsequence , in a given direction , for a specific layer :
Dominance score. To quantify the contributors to chat’s contextualization, we evaluate the attention sub-layer output in layer , by setting , in Eq. 2. can be any token subsequence preceding chat[-1] (e.g., adv).333Note that summing over the contributions of all tokens yields (i.e., ). To simplify the analysis to follow, we select the last token position before generation (chat[-1]) as a single representative token from chat. Formally:
Intuitively, this metric measures how much influences a target subsequence (chat[-1]), thus we refer to it as ’s dominance score. Mathematically, it captures the magnitude of ’s contribution in the direction of the total attention output at layer , which is in turn added to the residual stream (). When has a markedly higher dominance score than other token subsequences, we say it hijacks the context.
@@ 5.2 Hijacks
We use the dominance score (Eq. 5.1) to analyze how different subsequences contribute to the jailbreak-critical chat tokens across various prompt distributions. We find that suffix-based jailbreaks aggressively hijack chat and strongly suppress instr’s dominance, far more than other prompts.
We first compare two prompts with the same instruction—one using a random suffix (“!!..!”, failed sample) and the other a GCG suffix (successful)—and visualize their dominance scores across layers in Fig. 5. Unlike the random suffix, GCG dominates the chat representation from later layers (), in effect hijacking the contextualization process, while the instr’s influence gradually vanishes already from early layers (). Observing that this gap concentrates at the upper layers, we next analyze these.
Experimental setting. We extend our analysis by averaging the dominance score across the upper half of the layers (which can be seen as accumulating the respective area in Fig. 5) and evaluating it on 30 harmful instructions, each paired with suffixes from nine distributions (examples in Tab. 3). These distributions include both handcrafted and optimized suffixes; most of them, similarly to GCG’s affirmation objective, aim to influence the initial tokens of the response:
@@
Optimized Benign Injection: Using GCG’s optimizer to craft suffixes that maximize the likelihood for a specific benign start in the response (e.g., a greeting);
@@
Adversarial Injection wei2023jailbrokendoesllmsafety: aggressively prompting an affirmative start and cooperation. When used with harmful instructions, this serves as a handcrafted jailbreak prompt wei2023jailbrokendoesllmsafety.
vii.
Optimized Adversarial Injection: Using GCG’s optimizer on benign instructions, to maximize the likelihood of an affirmative response prefix;
viii.
BEAST Suffixes sadasivan2024fastadversarialattackslanguage: 30 BEAST jailbreak suffixes of varying universality (App. A.1), with prompts grouped as: successful, failed with affirmation (Fail[sure]; e.g., “Sure, I’ll help. I cannot.”), and other failures (Fail[~sure]).444Other suffix distributions (i-vii) are not grouped by attack success, as they mostly fail (see Fig. 11).
ix.
GCG Suffixes zou2023gcg-universaltransferableadversarialattacks: 100 GCG jailbreak suffixes (§3.1) of varying universality, following the same prompt groups of BEAST’s.
Results. Analyzing the suffixes’ dominance scores (Fig. 6), we find through GCG and BEAST that suffix-based jailbreaks strongly suppress the instruction, while exhibiting irregular dominance in contextualization. Jailbreak suffixes contribute a magnitude of over compared to other suffix distributions, including handcrafted jailbreaks and even benign GCG-like suffixes optimized to manipulate the response; this suggests that merely optimizing for a response prefix is not sufficient to reproduce GCG’s strong hijacking behavior. We observe similar trends when aggregating across all the layers (Fig. 15, App. B.3). Furthermore, GCG’s hijacking remains unusual even compared to dominant suffixes in benign instructions (from AlpacaEval; Fig. 16, App. B.3), underscoring the distinctiveness of this mechanism.
In general, successful GCG jailbreaks require heavily suppressing instr’s contribution, with adv strongly hijacking chat by contributing a large magnitude to the former’s representation (Figs. 6(a)–6(b)); as seen in §4, this predominant and unique mechanism is also critical for jailbreaks. GCG’s dominance extends to the entire prompt (Fig. 14), suggesting it additionally suppresses the influence of template tokens in pre-chat (e.g., <bos>) and chat itself—a pattern also visible in Fig. 5. Moreover, handcrafted jailbreaks’ relatively strong hijacking (e.g., compared with random suffixes) may explain their effective use as suffix initializers in jailbreak optimizers (primarily, through replacing random initialization) liu2024autodan, as we also later demonstrate (§7.1).
@@ 6 Hijacking is Key for GCG Universality
Experimental setting. We next evaluate Gemma2. We additionally validate our results on Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-32B, and Llama3.1-8B, and defer detailed analysis on Qwen2.5-1.5B to App. B.4. For Gemma2, we sample 350 GCG suffixes of diverse universality, along with 30 harmful instructions. For validation on the rest of the models, we sample 30 GCG suffixes, along with 10 harmful instructions. For each suffix, we average its adv dominance score in layer across the different instructions, referring to this measure as the suffix’s hijacking strength, and comparing it to the suffix universality score (calculated on a larger set of instructions; §3.1). Notably, this calculation involves a single forward pass.
Results. Fig. 7 shows the Spearman correlation between universality and layer-wise hijacking strengths (including a summation over all the layers), using an instruction set (of size 10) disjoint from the evaluations to follow. We find that dominance in the initial and final layers does not correlate with universality, whereas hijacking strength in later-mid layers does. Notably, layers 18–21 yield the highest Spearman correlations; specifically, layer 20 achieves a moderate correlation schober2018correlation of , -value , and confidence interval of fieller1957tests, with similar values for the other 20 instructions used in further evaluation.
@@
Notably, similar trends hold when measuring hijacking strength using a much smaller instruction set, even a single random harmful instruction (i.e., simply comparing suffixes’ dominance scores under an instruction; Fig. 8(b)). Additionally, to control for the possible effect of jailbreak success, we repeat this analysis using only failed attack samples (in particular, instruction-suffix pairs that elicit model refusal). As shown in Fig. 8(c), the trend persists, indicating the internal hijacking mechanism varies across suffixes, even when all produce the same refusal outcome.
Additionally, to validate the main trend (as presented in Fig. 8(a)) on different models, we repeat this process for Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-32B, and Llama3.1-8B, analyzing their hijacking strength in layers 15, 21, 35, and 14, respectively. The results, shown in Fig. 9, reaffirm the link between hijacking strength and universality, yielding Spearman correlations () of 0.425, 0.620, 0.650, and 0.719, respectively.
Lastly, we demonstrate that results and correlations are similar with other variants of hijacking strength (Fig. 17, App. B.4), underscoring the role of the attention scores and that hijacking can be inspected along a few directions in the model’s activation space. Specifically, hijacking strength (in layer 20) can also be computed as: (i) a statistic aggregating top advchat attention scores (which directly scale the transformed vectors, and could control the hijacking strength; Eq. 1); (ii) dominance score w.r.t. a principal direction in GCG jailbreaks (replacing the sample-specific attention activations in Eq. 5.1), derived using a difference-in-means approach on failed and successful GCG samples arditiRefusalLanguageModels2024.
@@ 7 Practical Implications
To further understand the hijacking phenomenon’s impact, and demonstrate its practical benefit, we apply our insights on GCG jailbreak to develop both offensive and defensive strategies. First, we show it is possible to boost GCG’s universality by encouraging hijacking throughout the attack (§7.1). Second, we successfully mitigate suffix-based jailbreaks by suppressing hijacking during inference (§7.2). Third, we systematically identify suffix-based jailbreak prompts by efficiently inspecting for hijacking (§7.3).
@@ 7.1 Boosting GCG Universality With Hijacking Enhancement
We leverage our insights on the relationship between hijacking and universality (§6) to encourage hijacking during GCG’s optimization. We show that this method allows us to obtain universal suffixes with a reduced computational cost, further highlighting the link between hijacking and attack strength.
@@
Existing approaches. Universal suffix-based jailbreaks are typically crafted by running the original affirmation objective on multiple harmful instructions simultaneously (GCG-Mult; zou2023gcg-universaltransferableadversarialattacks). However, optimizing GCG across instructions significantly increases computational cost—matching that of separate single-instruction runs.
@@
Hijacking-based approach (GCG-Hij). We propose a modified objective that is optimized against a single instruction, thus preserving the computational efficiency of single-instruction GCG. Specifically, motivated by the fact that attention scores both scale transformed vectors magnitude (thus enhance hijacking; Eq. 1), and increase with universality in middle layers (§6), we define an attention-score-based proxy objective (), which is then added to GCG’s affirmation loss to form GCG-Hij’s loss ():
@@
Experimental setting. We select 10 random instructions for attacks (disjoint from evaluation), run GCG and GCG-Hij on each instruction separately, and optimize GCG-Mult on all 10 instructions simultaneously.555We select values by line search on a dev set (disjoint from evaluation): , , and for Gemma2, Qwen2.5-1.5B, and Llama3.1, respectively, generally finding effective. For layers, we set and . We repeat this for 3 random seeds. To test whether GCG-Hij is more likely to yield universal suffixes from single-instruction optimization, we compute the average universality (§3.1) on the 10 suffixes crafted with GCG and GCG-Hij (Avg. Univ.). Then, under a unified budget of 10 optimized instructions, we compare GCG and GCG-Hij to GCG-Mult, reporting whether either surpasses GCG-Mult (Win Over GCG-Mult) and the fraction of such wins (% of suffixes won).
GCG-Hij boosts universality. Results are reported in Tab. 1. First, GCG-Hij achieves superior average universality compared to GCG (1.1–5); it is more likely to generate highly universal suffixes while incurring the same computational cost as the original single-instruction GCG. Second, under the same computational budget as GCG-Mult, GCG-Hij consistently yields multiple universal suffixes that individually outperform the single suffix produced by the former. These results substantiate the link between hijacking and universality, demonstrating that enhancing the former boosts the latter.
Furthermore, App. B.5 explores another GCG variant (GCG-HotInit) motivated by our findings: replacing the default, arbitrary initialization with a handcrafted jailbreak (which exhibits strong hijacking; §5) without modifying the original objective. GCG-HotInit yields universality gains comparable to GCG-Hij (Fig. 19(a)) and demonstrates a significant hijacking advantage early in the optimization (Fig. 19(b)).
7.2 Mitigating Suffix-based Jailbreaks With Hijacking Suppression
@@
Through our analysis, we found that GCG adversarial suffixes (adv) hijack chat’s representation in an irregular and often extreme manner, particularly for universal suffixes (§5–6), and that this hijacking underlies attack effectiveness (§4). We therefore hypothesize that surgically suppressing this hijacking could defend against suffix-based jailbreaks with minimal effect on benign prompts. To test this, we introduce and evaluate a training-free framework for Hijacking Suppression.
@@
Starting with (a), we consider as candidates all transformed vectors departing from the user input tokens (i.e., the user prompt, excluding special tokens) to chat tokens, denoted advinput. Notably, for GCG prompts, suppressing a large portion of these vectors (i.e., advchat, see §4.1), eliminates the attack. We use this general superset (advinput) so the framework remains applicable to any prompt, including benign ones, without prior prompt knowledge.
@@
Next, for (b), we assign each vector in the superset a score, and select the top-1%. Specifically, we use the attention score () for each transformed vector (), as it mathematically scales the vector (Eq. 1) potentially amplifying hijacking strength, and empirically, higher top-1% scores correlate with GCG suffix universality (§6). While we prioritize simplicity, future work may explore scoring methods that better disentangle jailbreak from model utility, or that avoid the overhead of materializing attention matrices—which is bypassed by optimized attention kernels dao2022flashattentionfastmemoryefficientexact.
@@
Finally, for (c), we suppress the top 1% of transformed vectors by scaling their magnitude by :666The transformed vector update is applied before layer normalization and is equivalent to reducing the corresponding post-softmax attention score.
@@
Experimental setting. We apply Hij. Suppr. with on different models,777We found balances robustness and utility; further tuning may improve results. and evaluate the effect on: (i) model robustness, by measuring the attack success rate on challenging custom datasets of 1.5K GCG jailbreak prompts and 1.5K BEAST jailbreak prompts, composed of harmful instructions from AdvBench zou2023gcg-universaltransferableadversarialattacks and StrongReject souly2024strongrejectjailbreaks, each appended to various jailbreak suffixes, that originally led to diverse attack success; and on (ii) model utility, using AlpacaEval alpaca-eval2023 and MMLU mmlu-hendrycks2021measuringmassivemultitasklanguage, common evaluations for LLM helpfulness llm-eval-survey-chang-2024. See App. A.3 for more technical details.
| Attack Success () | Utility () | ||||
| GCG | BEAST | AlpacaEval | MMLU | ||
| Gemma2 | Initial | ||||
| +Hij. Suppr. | |||||
| Qwen2.5-1.5B | Initial | ||||
| +Hij. Suppr. | |||||
| Llama3.1 | Initial | ||||
| +Hij. Suppr. | |||||
@@
Hij. Suppr. improves robustness vs. jailbreaks. Tab. 2 presents Hij. Suppr.’s effect on attack success. Specifically, it substantially mitigates suffix-based jailbreaks, reducing success rates by a factor of –. Crucially, we observe only a marginal decrease in model utility, with drops of 2% on MMLU and AlpacaEval. AlpacaEval responses remain highly similar after applying Hij. Suppr., with average RougeL scores of –, indicating minimal change lin-2004-rouge. Still, we expect further refinement of the framework (e.g., the scoring step, (b)) to improve robustness-utility tradeoff.
7.3 Detecting Suffix-based Jailbreaks Through Hijacking
Our analysis shows that jailbreak suffixes leave a distinct signature: hijacking of the chat’s contextualization. We hypothesize this can also be leveraged to reliably distinguish malicious prompts from benign ones, prior to the generation of the response. We demonstrate the feasibility of such a detection scheme, further highlighting the distinct hijacking of suffix-based jailbreaks.
Hijacking-based Detection (Hij. Detect.). Our scheme consists of (1) an offline step of attention-head selection and (2) an online step of classification, inspired by the Prompt Injection classifier proposed by hung-etal-2025-attention.
First, we identify the attention heads most indicative of jailbreak suffixes’ hijacking. We do this by analyzing a training set of 100 GCG jailbreak prompts and 100 benign instructions (AlpcaEval; alpaca-eval2023): we measure the average dominance score (from all user input tokens to chat) for both prompt distributions, and select the top five (2%) attention heads with the largest score gap between the two distributions.888Testing different numbers of top heads on the training set, we find selecting 2%–15% to perform similarly. Then, after picking the critical heads, we classify prompts as follows: given a user input, compute the dominance score (inputchat) summarized only on the selected heads (requires recording only a single forward pass), and use this as the detection score—i.e., high scores indicate attacks.
Experimental setting. We evaluate Hij. Detect. for detecting a positive set of successful jailbreak prompts from GCG and BEAST (1K prompts each; disjoint from the train set), against a negative set of non-jailbreak prompts (i.e., resulting in benign responses). The negative set is composed of: 1K benign instructions (from AlpacaEval and HuggingFace Instruction dataset;999https://huggingface.co/datasets/HuggingFaceH4/instruction-dataset disjoint from the train set) and 0.5K non-jailbreak harmful prompts (instructions from AdvBench, without jailbreak attempts). We use the mere harmful instructions as negatives to verify that our method detects the jailbreak itself, rather than general input harmfulness.
Hij. Detect. detects suffix-based jailbreaks. Results in Fig. 10, show that the hijacking-based classifier achieves a Receiver Operating Characteristic (ROC) Area Under Curve (AUC) of 0.95, effectively distinguishing jailbreak and non-jailbreak prompts. This demonstrates a potential for leveraging hijacking for efficient jailbreak detection.